A caterer charges $120 to cater a party for 15 people
and $200 for 25 people. Assume that the cost, y, is a
linear function of the number of x people. Write an
equation in slope-intercept form for this function.
What does the slope represent? How much would a
party for 40 people cost?
y = 15x + 120
y = 8x + 120
y = 8x
y = 10x + 80

Answer by greenestamps(13198)



The cost goes up by $200-$120 = $80 when the number of people goes up by 25-15 = 10; the slope is 80/10 = 8. So the equation is of the form

y = 8x+b

Use either given data point to find b:

120 = 8(15)+b
120 = 120+b
b = 0

ANSWER: y = 8x

The slope of 8 means the cost goes up by $8 for each additional person.

The cost for a party for 40 people is 40($8) = $320.